# Requirements Coverage-Guided Minimization for Natural Language Test Cases

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: xxx; 2025; XXXXisbn: 978-1-4503-XXXX-X/2018/06copyright: ccjournal: TOSEMjournalyear: 2026journalvolume: 1journalnumber: 1publicationmonth: 1doi: 10.1145/3798164ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Requirements analysis
@@ 2.1. Test Cases based on Automotive Requirements
$2.$ Test Suite Minimization
@@ 3. Approach
This section introduces RTM, a TSM approach that satisfies the three criteria outlined in Section $2.$ It targets test cases derived from system requirements for validation. In short, it leverages a text embedding technique combined with a distance function to compute test case similarity, then employs evolutionary search to minimize the test suite. The primary objective of RTM is to find the optimal $\mathcal{T}^{\prime}$ that maximizes fault detection capability by minimizing the similarity values between the test cases while reducing the testing cost, measured in terms of test suite size, as well as ensuring $100\%$ requirement coverage.
@@ 3.2. Test Case Representation
In order to bound the WMD between 0 and 11, and convert it from a distance function to a similarity measure, we normalize it as follows:
@@ 4.2. Dataset
For example, for a test suite $\mathcal{T}$ with three test cases that detect 33, 22, and 33 faults, respectively. The total number of unique faults detected by $\mathcal{T}$ is 44, the redundancy level of $\mathcal{T}$ is therefore $(3+2+3)/4=2$. Similarly, the redundancy level of our dataset is $2610/220=11.86$.
@@ 4.3. Text Representations
Word2Vec (Mikolov et al., 2013). Word2Vec has been shown to be effective for identifying similar test cases (Viggiato et al., 2022). In this study, we used the Continuous Bag-of-Words (CBOW) architecture with a window size of $10$ for Word2Vec, as this configuration yielded the highest FDR for TSM. This was determined through experiments evaluating both the CBOW and Skip-Gram architectures with window sizes ranging from 22 to $10$, in increments of 22. We trained Word2Vec models on the entire test case corpus and transformed each word into $300$-dimensional numerical vectors.
GloVe (Pennington et al., 2014). GloVe is an unsupervised learning algorithm that leverages global word-word co-occurrence statistics from a corpus to learn word embeddings (Pennington et al., 2014). Its core idea is that the ratio of co-occurrence probabilities can capture meaningful linear substructures within the word vector space. We selected a window size of 22 for GloVe based on its performance in hyperparameter tuning experiments with window sizes ranging from 22 to $10$, in increments of 22. As in Word2Vec, we used GloVe to generate a $300$-dimensional numerical vector representation for each word in the test cases.
FastText (Bojanowski et al., 2017). FastText is a word embedding model proposed by Bojanowski et al. (Bojanowski et al., 2017), which extends Word2Vec by incorporating subword information. Each word is represented as a combination of character n-grams, enabling the model to better handle rare and out-of-vocabulary words, and to generalize more effectively—particularly in morphologically rich or noisy text. We used the CBOW architecture with a window size of 88 as the FastText configuration, based on experiments evaluating both CBOW and Skip-Gram architectures with window sizes ranging from 22 to $10$ in increments of 22. We used FastText to convert each word in the test cases into a $300$-dimensional numerical vector.
$4.$ Baselines
@@ 5.2. Performance of RTM compared with baselines (RQ2)
We observe that RTM consistently outperforms Random Minimization and FAST-R approaches in terms of FDR across various minimization budgets. The performance gap is particularly notable at mid-range budgets ($30\%$ to $60\%$). For instance, at a $40\%$ minimization budget, RTM achieves an FDR of $78.68\%$, whereas RM-Req and FAST++ yield approximately $65.45\%$ and $56.50\%$, respectively. This demonstrates RTM’s superior ability to retain fault-revealing test cases, especially under mid-range budget constraints. At lower and higher minimization budgets, the FDR of all techniques tends to converge toward either 0 or 11, thus making the difference in FDR between RTM and other techniques smaller. However, mid-range minimization budgets are more important in practice, as they offer more useful trade-offs between cost savings and fault detection.
@@ 5.3. Impact of test suites redundancy (RQ3)
As expected, the FDR of all techniques improves as the test suite redundancy level increases, indicating that the TSM techniques tend to perform better–achieving higher FDR–when applied to more redundant test suites. This trend, however, becomes less evident at higher budgets ($80\%$ and $90\%$), where the FDR of all techniques converge toward 11. Random Minimization (RM-Req and RM-NoReq) slightly outperforms other FAST-R variants but consistently falls short of RTM. We also observed that across all minimization budgets and redundancy levels, RM-Req consistently outperformed RM-NoReq in FDR. This demonstrates that, for our dataset, explicitly ensuring that every requirement is covered not only preserves full requirement coverage but also significantly boosts the test suite’s ability to detect faults. In other words, prioritizing requirement coverage is an effective heuristic for improving fault detection under budgetary constraints.
@@ 5.4. Scalability of RTM compared with baselines (RQ4)
In summary, RTM runs within 55 seconds across all evaluated test suites. Although being slower than the baselines, RTM scales approximately linearly with test suite sizes, and even for substantially larger suites, the absolute runtime gaps remain acceptable in practice—they are not expected to have practical consequences for TSM. Given that RTM consistently achieves much higher FDR across various datasets and minimization budgets than the baselines while preserving full requirement coverage, it is therefore a better option in many practical contexts.
@@ 7. Related Work
Greedy-based approaches follow a heuristic, step-by-step selection strategy, where the test case that provides the most benefit toward the objective is chosen at each iteration, until the given constraints (e.g., full requirement coverage or code path coverage) are satisfied. Miranda et al. (Miranda and Bertolino, 2017) and Noemmer et al. (Noemmer and Haas, 2019) both employed greedy heuristics for TSM based on statement coverage. Their approaches iteratively selected test cases to maximize code coverage, achieving reasonable FDRs while significantly reducing the number of executed test cases. Greedy-based approaches are computationally efficient and easy to implement, but may suffer from suboptimal global performance due to their locally optimal decisions. Moreover, given that, in our dataset, each test case covers only one requirement, the coverage-based greedy strategy–iteratively selecting the test cases that maximize the coverage until the full coverage constraint is satisfied–is equivalent to Random Minimization with the requirement constraint (RM-Req), as described in Section $4.$

### v4

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3798164isbn: 978-1-4503-XXXX-X/2018/06copyright: ccjournal: TOSEMjournalyear: 2026journalvolume: 1journalnumber: 1publicationmonth: 1doi: 10.1145/3798164ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Requirements analysis
@@ 2.1. Test Cases based on Automotive Requirements
2.2. Test Suite Minimization
@@ 3. Approach
This section introduces RTM, a TSM approach that satisfies the three criteria outlined in Section 2.2. It targets test cases derived from system requirements for validation. In short, it leverages a text embedding technique combined with a distance function to compute test case similarity, then employs evolutionary search to minimize the test suite. The primary objective of RTM is to find the optimal $\mathcal{T}^{\prime}$ that maximizes fault detection capability by minimizing the similarity values between the test cases while reducing the testing cost, measured in terms of test suite size, as well as ensuring $100\%$ requirement coverage.
@@ 3.3. Similarity Measure
In order to bound the WMD between 0 and $1$, and convert it from a distance function to a similarity measure, we normalize it as follows:
@@ 4.2. Dataset
For example, for a test suite $\mathcal{T}$ with three test cases that detect $3$, 22, and $3$ faults, respectively. The total number of unique faults detected by $\mathcal{T}$ is $4$, the redundancy level of $\mathcal{T}$ is therefore $(3+2+3)/4=2$. Similarly, the redundancy level of our dataset is $2610/220=11.86$.
@@ 4.3. Text Representations
Word2Vec (Mikolov et al., 2013). Word2Vec has been shown to be effective for identifying similar test cases (Viggiato et al., 2022). In this study, we used the Continuous Bag-of-Words (CBOW) architecture with a window size of $10$ for Word2Vec, as this configuration yielded the highest FDR for TSM. This was determined through experiments evaluating both the CBOW and Skip-Gram architectures with window sizes ranging from $2$ to $10$, in increments of $2$. We trained Word2Vec models on the entire test case corpus and transformed each word into $300$-dimensional numerical vectors.
GloVe (Pennington et al., 2014). GloVe is an unsupervised learning algorithm that leverages global word-word co-occurrence statistics from a corpus to learn word embeddings (Pennington et al., 2014). Its core idea is that the ratio of co-occurrence probabilities can capture meaningful linear substructures within the word vector space. We selected a window size of $2$ for GloVe based on its performance in hyperparameter tuning experiments with window sizes ranging from $2$ to $10$, in increments of $2$. As in Word2Vec, we used GloVe to generate a $300$-dimensional numerical vector representation for each word in the test cases.
FastText (Bojanowski et al., 2017). FastText is a word embedding model proposed by Bojanowski et al. (Bojanowski et al., 2017), which extends Word2Vec by incorporating subword information. Each word is represented as a combination of character n-grams, enabling the model to better handle rare and out-of-vocabulary words, and to generalize more effectively—particularly in morphologically rich or noisy text. We used the CBOW architecture with a window size of $8$ as the FastText configuration, based on experiments evaluating both CBOW and Skip-Gram architectures with window sizes ranging from $2$ to $10$ in increments of $2$. We used FastText to convert each word in the test cases into a $300$-dimensional numerical vector.
4.4. Baselines
@@ 5.2. Performance of RTM compared with baselines (RQ2)
We observe that RTM consistently outperforms Random Minimization and FAST-R approaches in terms of FDR across various minimization budgets. The performance gap is particularly notable at mid-range budgets ($30\%$ to $60\%$). For instance, at a $40\%$ minimization budget, RTM achieves an FDR of $78.68\%$, whereas RM-Req and FAST++ yield approximately $65.45\%$ and $56.50\%$, respectively. This demonstrates RTM’s superior ability to retain fault-revealing test cases, especially under mid-range budget constraints. At lower and higher minimization budgets, the FDR of all techniques tends to converge toward either 0 or $1$, thus making the difference in FDR between RTM and other techniques smaller. However, mid-range minimization budgets are more important in practice, as they offer more useful trade-offs between cost savings and fault detection.
@@ 5.3. Impact of test suites redundancy (RQ3)
As expected, the FDR of all techniques improves as the test suite redundancy level increases, indicating that the TSM techniques tend to perform better–achieving higher FDR–when applied to more redundant test suites. This trend, however, becomes less evident at higher budgets ($80\%$ and $90\%$), where the FDR of all techniques converge toward $1$. Random Minimization (RM-Req and RM-NoReq) slightly outperforms other FAST-R variants but consistently falls short of RTM. We also observed that across all minimization budgets and redundancy levels, RM-Req consistently outperformed RM-NoReq in FDR. This demonstrates that, for our dataset, explicitly ensuring that every requirement is covered not only preserves full requirement coverage but also significantly boosts the test suite’s ability to detect faults. In other words, prioritizing requirement coverage is an effective heuristic for improving fault detection under budgetary constraints.
@@ 5.4. Scalability of RTM compared with baselines (RQ4)
In summary, RTM runs within $5$ seconds across all evaluated test suites. Although being slower than the baselines, RTM scales approximately linearly with test suite sizes, and even for substantially larger suites, the absolute runtime gaps remain acceptable in practice—they are not expected to have practical consequences for TSM. Given that RTM consistently achieves much higher FDR across various datasets and minimization budgets than the baselines while preserving full requirement coverage, it is therefore a better option in many practical contexts.
@@ 7. Related Work
Greedy-based approaches follow a heuristic, step-by-step selection strategy, where the test case that provides the most benefit toward the objective is chosen at each iteration, until the given constraints (e.g., full requirement coverage or code path coverage) are satisfied. Miranda et al. (Miranda and Bertolino, 2017) and Noemmer et al. (Noemmer and Haas, 2019) both employed greedy heuristics for TSM based on statement coverage. Their approaches iteratively selected test cases to maximize code coverage, achieving reasonable FDRs while significantly reducing the number of executed test cases. Greedy-based approaches are computationally efficient and easy to implement, but may suffer from suboptimal global performance due to their locally optimal decisions. Moreover, given that, in our dataset, each test case covers only one requirement, the coverage-based greedy strategy–iteratively selecting the test cases that maximize the coverage until the full coverage constraint is satisfied–is equivalent to Random Minimization with the requirement constraint (RM-Req), as described in Section 4.4.